# Semiclassics for the QCD vacuum structure through $T^{2}$-compactification with the baryon-’t Hooft flux

### v1

@@ 5.2 Comments on domain wall
Note that the domain wall becomes non-dynamical in the large mass limit $m\rightarrow\infty$. The absence of a domain-wall phase transition is attributed to the problem that the domain wall is non-dynamical in $SU(N)$ Yang-Mills theory in our framework. In the original 4d $SU(N)$ Yang-Mills theory, the domain wall is a dynamical object described by the $SU(N)$ level-$1$ Chern-Simons theory. However, in our semiclassical treatment for $SU(N)$ Yang-Mills theory, the domain wall between two distinct vacuum labels is non-dynamical [^12] As explained in Gaiotto:2017tne , in the original 4d theory, it becomes favorable for the $\eta^{\prime}$ field to “jump over the singular point” (which means the shift of the vacuum label) at a certain value of mass rather than to interpolate the two minima smoothly. In the semiclassical treatment, the “jumping over the singular point” of the $\eta^{\prime}$ potential is energetically unfavorable because the gluonic fields are almost frozen at the small $T^{2}$ compactification..
@@ 6.1 $CP$ breaking at $\theta=\pi$
This ansatz reduces the variables to ($\varphi,k$) $\in U(1)\times\mathbb{Z}_{N}$. Let us take the branch of $\varphi$ to be $-\pi<\varphi<\pi$, and then the potential $V[\eta^{\prime},\varphi_{1},\cdots,\varphi_{N_{f}-1}]$ becomes
@@ 6.2 Comments on baryon-color-flavor anomalies
Previously, within the ansatz, the semiclassical description consists of reduced variables $(\varphi,k)\in U(1)\times\mathbb{Z}_{N}$. Notably, when $\operatorname{gcd}(N,N_{f})=1$ and only in this case, the label $k$ can be absorbed by extending $\varphi$: $(N_{f}\varphi+2\pi k)\rightarrow N_{f}\varphi~{}~{}(\operatorname{mod}2\pi N)$. Thus, we can express the periodic scalar $\varphi\sim\varphi+2\pi N$ with the potential
@@
with arbitrary integers $K_{1},K_{2}$ are allowed as symmetric perturbations. When $\operatorname{gcd}(N,N_{f})=1$, we can choose integers $K_{1}$ and $K_{2}$ such that $NK_{1}-N_{f}K_{2}=1$, which ends up with giving the $\cos\left(\frac{\varphi-\theta}{N}\right)$ term. Therefore, if $\operatorname{gcd}(N,N_{f})=1$, we can achieve a trivial vacuum for all $\theta\sim\theta+2\pi$ by perturbation of symmetry-allowed operators, which means the absence of anomaly or global inconsistency.
@@
As the 2d effective theory is derived from the assumption of adiabatic continuity, it is less significant to discuss quantitative aspects. Nevertheless, as this semiclassical description naturally bridges the chiral limit and the quenched limit, it would be worthwhile to study some quantities related to the vacuum structure. In the rest of this section, let us make some observations of the pion mass and topological susceptibility. Here, we estimate the pion mass at a small quark mass. Its asymptotic behavior in the massless limit could be possibly kept by adiabatic continuity.
@@
where $U$ is redefined as $U\rightarrow\mathrm{e}^{\mathrm{i}\varphi_{*}}U$ to denote the deviation from the minimum $\mathrm{e}^{\mathrm{i}\varphi_{*}}\textbf{1}$. Thus, the “pion” mass can be estimated as $m^{2}_{\pi}\sim m\mu\cos(\varphi_{*})$. For general $(N_{f},|\theta|<\pi)$, it is $m^{2}_{\pi}=O(m)$ in accordance with the orthodox PCAC relation.
@@ 6.4 Topological susceptibility
One can extract the topological susceptibility from the $\theta$ dependence of the vacuum energy: $\mathcal{E}=\frac{1}{2}\chi_{\mathrm{top}}\theta^{2}+O(\theta^{4})$. From the perturbation theory in $\theta$, we have,
@@ 7.2 Dashen phase of $(1+1+1)$-flavor QCD
The above phase diagram may be affected by the strange quark. Next, let us consider the phase diagram of $N_{f}=1+1+1$ QCD on the $(m_{u},m_{d})$ plane with fixed strange mass $m_{s}>0$ and $\theta=0$. For large $m_{s}$, the qualitative feature of the phase diagram is unchanged as it is essentially $(1+1)$-flavor QCD. A nontrivial phenomenon happens if the strange mass is below the critical mass of the 1-flavor QCD (see Fig. 3). In this case, the $CP$ restored phase appears in the limit $(m_{u},m_{d})\to(\infty,-\infty)$ and $(m_{u},m_{d})\to(-\infty,\infty)$.
@@
  • In Creutz:2003xu , the $U(3)$ chiral Lagrangian with $(\operatorname{det}U)$-type $\eta^{\prime}$ mass term was also examined, concluding that it does not qualitatively differ from the chiral Lagrangian. This incorporation of the $\eta^{\prime}$ meson gives the qualitatively same phase diagram as Fig. 5, However, even with a heavy strange mass $m_{s}>\Delta$, the phases (C) and (D) artificially persist in the $U(3)$ chiral Lagrangian [^20] For instance, instead of phase (C), the $U(3)$ chiral Lagrangian yields an artificial trivial phase for $m_{s}>\Delta$: At this vacuum, the chiral field $U$ is, $U=\begin{pmatrix}1&0&0\\ 0&-1&0\\ 0&0&1\end{pmatrix}$ which is equivalent to $\varphi_{1}=0,\varphi_{2}=\pi,\eta^{\prime}=\pi$ in our parameterization. In our model, this configuration is not $CP$ invariant: $(\varphi_{1},\varphi_{2},\eta^{\prime})=(0,\pi,\pi)$ and $(\varphi_{1},\varphi_{2},\eta^{\prime})=(0,\pi,-\pi)$ are discriminated due to the periodicity extension (See also footnote 18). . As previously discussed, this artificial phase is inconsistent with the phase diagram of the one-flavor QCD at $\theta=\pi$ (Fig. 3) in the decoupling limit of $m_{u}\rightarrow+\infty$ and $m_{d}\rightarrow-\infty$. The use of $(2\pi N)$-periodic $\eta^{\prime}$ meson improves the consistency with global aspects.
@@ Acknowledgements.
The authors thank Hiromasa Watanabe for useful discussion. The work of Y. T. was supported by Japan Society for the Promotion of Science (JSPS) KAKENHI Grant numbers, 22H01218, and by Center for Gravitational Physics and Quantum Information (CGPQI) at Yukawa Institute for Theoretical Physics. Y. H. was supported by JSPS Research Fellowship for Young Scientists Grant No. 23KJ1161
@@ Appendix A Bosonized Schwinger model: an exercise for extending periodicity
A common manipulation is to integrate out the $U(1)$ gauge field, and we get the mass term $\frac{e^{2}}{8\pi^{2}}(\phi+\theta)^{2}$. Naively, this appears inconsistent with the periodicity. Now, let us carefully consider the integration of the $U(1)$ gauge field step-by-step.
@@
and its integration [^21] If the summation over the constant holonomy were over $\mathbb{Z}_{N}$, the constrain would be $c/(2\pi N)\in H^{1}(M_{2};\mathbb{Z})$. This happens in the main text, Sec. 3.2.2. constraints the nontrivial winding: $c/2\pi=0$. This enables us to regard $\phi$ as a globally-defined $\mathbb{R}$-valued field.
@@
We note that $\phi$ in the last expression is subject to the constraint $[\mathrm{d}\phi]=0\in H^{1}(M_{2};\mathbb{Z})$, which means that $\phi$ cannot have a nontrivial winding, and this constraint appears as a result of the integration over the holonomies.

### v3

@@ 5.2 Comments on domain wall
Note that the domain wall becomes non-dynamical in the large mass limit $m\rightarrow\infty$. The absence of a domain-wall phase transition is attributed to the problem that the domain wall is non-dynamical in $SU(N)$ Yang-Mills theory in our framework. In the original 4d $SU(N)$ Yang-Mills theory, the domain wall is a dynamical object described by the $SU(N)$ level-$1$ Chern-Simons theory. However, in our semiclassical treatment for $SU(N)$ Yang-Mills theory, the domain wall between two distinct vacuum labels is non-dynamical [^18] As explained in Gaiotto:2017tne , in the original 4d theory, it becomes favorable for the $\eta^{\prime}$ field to “jump over the singular point” (which means the shift of the vacuum label) at a certain value of mass rather than to interpolate the two minima smoothly. In the semiclassical treatment, the “jumping over the singular point” of the $\eta^{\prime}$ potential is energetically unfavorable because the gluonic fields are almost frozen at the small $T^{2}$ compactification (see footnote 10 and also the last two paragraphs of Sec. 2 on the domain wall)..
@@ 6.1 $CP$ breaking at $\theta=\pi$
This ansatz reduces the variables to ($\mathrm{e}^{\mathrm{i}\varphi},k$) $\in U(1)\times\mathbb{Z}_{N}$. Let us take the branch of $\varphi$ to be $-\pi<\varphi<\pi$, and then the potential $V[\eta^{\prime},\varphi_{1},\cdots,\varphi_{N_{f}-1}]$ becomes
@@ 6.2 Comments on baryon-color-flavor anomalies
Previously, within the ansatz, the semiclassical description consists of reduced variables $(\mathrm{e}^{\mathrm{i}\varphi},k)\in U(1)\times\mathbb{Z}_{N}$. Notably, when $\operatorname{gcd}(N,N_{f})=1$ and only in this case, the label $k$ can be absorbed by extending $\varphi$: $(N_{f}\varphi+2\pi k)\rightarrow N_{f}\varphi~{}~{}(\operatorname{mod}2\pi N)$. Thus, we can express the periodic scalar $\varphi\sim\varphi+2\pi N$ with the potential
@@
with arbitrary integers $K_{1},K_{2}$ are allowed as symmetric perturbations. When $\operatorname{gcd}(N,N_{f})=1$, we can choose integers $K_{1}$ and $K_{2}$ such that $NK_{1}-N_{f}K_{2}=1$, which ends up with giving the $\cos\left(\frac{\varphi-K_{2}\theta}{N}\right)$ term. Therefore, if $\operatorname{gcd}(N,N_{f})=1$, we can achieve a trivial vacuum for all $\theta\sim\theta+2\pi$ by perturbation of symmetry-allowed operators, which means the absence of anomaly or global inconsistency.
@@
As the 2d effective theory is derived from the assumption of adiabatic continuity, it is less significant to discuss quantitative aspects. Nevertheless, as this semiclassical description naturally bridges the chiral limit ($m\rightarrow 0$) and the quenched limit ($m\rightarrow\infty$), it would be worthwhile to study some quantities related to the vacuum structure. In the rest of this section, let us make some observations of the pion mass and topological susceptibility. Here, we estimate the pion mass at a small quark mass. Its asymptotic behavior in the massless limit could be possibly kept by adiabatic continuity.
@@
where $U$ is redefined as $U\rightarrow\mathrm{e}^{\mathrm{i}\varphi_{*}}U$ to denote the deviation from the minimum $\mathrm{e}^{\mathrm{i}\varphi_{*}}\textbf{1}$. Thus, the “pion” mass can be estimated as $m^{2}_{\pi}\sim m\mu\cos(\varphi_{*})$. For general $(N_{f},|\theta|<\pi)$, it is $m^{2}_{\pi}=O(m)$ in accordance with the orthodox partially conserved axial current (PCAC) relation.
@@ 6.4 Topological susceptibility
The topological susceptibility $\chi_{\mathrm{top}}$ is defined as the correlation of the topological charge density. This quantity is often studied as a characteristic quantity that represents the fluctuation of the topological charge, e.g., see Vicari:2008jw for a review. In terms of the vacuum energy, it quantifies the response to changes in the $\theta$ angle: $\mathcal{E}=\frac{1}{2}\chi_{\mathrm{top}}\theta^{2}+O(\theta^{4})$.
We can obtain the topological susceptibility from the perturbation theory in $\theta$ for the vacuum energy:
@@ 7.2 Dashen phase of $(1+1+1)$-flavor QCD
The above phase diagram is affected by the strange quark. Next, let us consider the phase diagram of $N_{f}=1+1+1$ QCD on the $(m_{u},m_{d})$ plane with fixed strange mass $m_{s}>0$ and $\theta=0$. For large $m_{s}$, the qualitative feature of the phase diagram is unchanged as it is essentially $(1+1)$-flavor QCD. A nontrivial phenomenon happens if the strange mass is below the critical mass of the 1-flavor QCD (see Fig. 3). In this case, the $CP$ restored phase appears in the limit $(m_{u},m_{d})\to(\infty,-\infty)$ and $(m_{u},m_{d})\to(-\infty,\infty)$.
@@
  • In Creutz:2003xu , the $U(3)$ chiral Lagrangian with $(\operatorname{det}U)$-type $\eta^{\prime}$ mass term was also examined, concluding that it does not qualitatively differ from the chiral Lagrangian. This incorporation of the $\eta^{\prime}$ meson gives the qualitatively same phase diagram as Fig. 5, However, even with a heavy strange mass $m_{s}>\Delta$, the phases (C) and (D) artificially persist in the $U(3)$ chiral Lagrangian [^26] For instance, instead of phase (C), the $U(3)$ chiral Lagrangian yields an artificial trivial phase for $m_{s}>\Delta$: At this vacuum, the chiral field $U$ is, $U=\begin{pmatrix}1&0&0\\ 0&-1&0\\ 0&0&1\end{pmatrix}$ which is equivalent to $\varphi_{1}=0,\varphi_{2}=\pi,\eta^{\prime}=\pi$ in our parameterization. In our model, this configuration is not $CP$ invariant: $(\varphi_{1},\varphi_{2},\eta^{\prime})=(0,\pi,\pi)$ and $(\varphi_{1},\varphi_{2},\eta^{\prime})=(0,\pi,-\pi)$ are discriminated due to the periodicity extension (See also footnote 24). . As previously discussed, this artificial phase is inconsistent with the phase diagram of the one-flavor QCD at $\theta=\pi$ (Fig. 3) in the decoupling limit of $m_{u}\rightarrow+\infty$ and $m_{d}\rightarrow-\infty$. The use of $(2\pi N)$-periodic $\eta^{\prime}$ meson improves the consistency with global aspects.
@@ Acknowledgements.
The authors thank Sinya Aoki, Aleksey Cherman, and Hiromasa Watanabe for useful discussions and comments on the draft. The work of Y. T. was supported by Japan Society for the Promotion of Science (JSPS) KAKENHI Grant numbers, 22H01218, and by Center for Gravitational Physics and Quantum Information (CGPQI) at Yukawa Institute for Theoretical Physics. Y. H. was supported by JSPS Research Fellowship for Young Scientists Grant No. 23KJ1161
@@ Appendix A Bosonized Schwinger model: an exercise for extending periodicity
A common manipulation is to integrate out the $U(1)$ gauge field, and we get the mass term $\frac{e^{2}}{8\pi^{2}}(\phi+\theta)^{2}$. Naively, this appears inconsistent with the periodicity, and let us carefully consider the integration of the $U(1)$ gauge field step-by-step (see also Armoni:2018bga ; Misumi:2019dwq ; Tanizaki:2019rbk ; Cherman:2020cvw ; Cherman:2021nox ; Cherman:2022ecu ).
@@
and its integration [^27] If the summation over the constant holonomy were over $\mathbb{Z}_{N}$, the constrain would be $c/(2\pi N)\in H^{1}(M_{2};\mathbb{Z})$. This happens in the main text, Sec. 3.2.2. constraints the nontrivial winding: $c/2\pi=0$. For example, on $T^{2}$ with the size $L$, the $2\pi$-periodic field $\phi$ may have the globally winding configuration such as $\phi(x_{1},x_{2})=2\pi(n_{1}\frac{x_{1}}{L}+n_{2}\frac{x_{2}}{L})$ with $n_{1},n_{2}\in\mathbb{Z}$, but the holonomy integration sets $n_{1},n_{2}=0$. This enables us to regard $\phi$ as a globally-defined $\mathbb{R}$-valued field.
@@
We note that $\phi$ in the last expression is subject to the constraint $[\mathrm{d}\phi]=0\in H^{1}(M_{2};\mathbb{Z})$, which means that $\phi$ cannot have a nontrivial global winding, and this constraint appears as a result of the integration over the holonomies.